# Spectrum Sharing in Satellite-Terrestrial Integrated Networks: Frameworks, Approaches, and Opportunities

### v2

@@ Abstract
With the construction of low-earth orbit (LEO) satellite constellations, ubiquitous connectivity has been achieved. However, to accommodate the increasing communication needs in non-terrestrial networks (NTNs), wireless users in remote areas may require access to more spectrum than is currently allocated. Terrestrial networks (TNs), such as cellular networks, are deployed in specific areas, but many underused licensed spectrum bands remain in remote areas. Therefore, bringing NTNs to a shared spectrum with TNs can improve network capacity under reasonable interference management. However, in satellite-terrestrial integrated networks (STINs), the comprehensive coverage of a satellite and the unbalanced communication resources of STINs make it challenging to manage mutual interference between NTN and TN effectively. This article presents the fundamentals and prospects of spectrum sharing (SS) in STINs by introducing four SS frameworks, their potential application scenarios, and technical challenges. Furthermore, advanced SS approaches related to interference management in STINs and performance metrics of SS in STINs are introduced. Moreover, a preliminary performance evaluation showcases the potential for sharing the spectrum between NTN and TN. Finally, future research opportunities for SS in STINs are discussed.
@@ I Introduction
Due to the boosting need for broader and more reliable wireless network coverage and data transmission, non-terrestrial networks (NTNs) have emerged as a promising solution. In urban areas with heavy traffic or remote areas where terrestrial networks (TNs) face geographical deployment limitations, NTNs can provide seamless access and ubiquitous connectivity [1]. Spectrum integration in satellite-terrestrial integrated networks (STINs) is a necessary step to improve the utilization efficiency of scarce spectrum resources. However, the coexistence of NTNs and TNs may be challenging, especially with regard to spectrum utilization and interference management. When NTN nodes such as low-earth orbit (LEO) satellites are incorporated into existing TNs via certain frequency bands, interference patterns can become more complicated due to different spatial propagation characteristics of signals. For example, the use of the Ka band or the Ku band for satellites may cause severe co-channel interference in existing communication systems. Moreover, within a fixed spectrum allocation, it is possible that while certain portions of the spectrum in NTNs are fully occupied, other segments within TNs may remain unutilized. The imbalance will lead to under-utilization of available spectrum resources and a significant hindrance to substantial improvements in system performance.
@@
In recent years, the rapid growth of NTNs has paved the way toward realizing STINs. STINs, a promising candidate for future sixth-generation (6G) wireless communications, incorporate LEO nodes into existing TN to take advantage of NTN and meet diverse communication requirements [2]. By using STINs, the coverage, flexibility, and spectrum efficiency of the wireless network can be significantly promoted, as terrestrial users can access various TN and NTN nodes from vast geographical areas. Inevitably, the spectrum resource will become scarcer in STIN as more NTN nodes are introduced. This encourages us to manage the spectrum when TN and NTN co-exist in the STIN [3]. Meanwhile, the licensed spectrum to the static TN nodes cannot always be utilized efficiently, as spatial spectrum holes cannot be eliminated within the geographic area of interest. Hence, mobile NTN users can admit access to users within the spectrum holes and promote spectrum efficiency [4]. In existing work, one of the promising technologies for TN and NTN coexistence is spectrum sharing (SS), which allows the NTN to share the spectrum pre-allocated to the TN. SS allows the NTN to work in the same frequency band when the TN users are active, as long as its interference to the TN can be tolerated [4, 5]. In this way, the STIN can provide high-quality and seamless wireless services at a slight cost of performance loss under the practical restrictions of limited spectrum resources.
@@
With the distinct advantages above, STINs can overwhelm conventional TNs in a variety of practical fields. As the key technique in STINs, SS becomes the main bottleneck in network performance promotion and faces several challenges [6]. First, a universal SS architecture cannot always meet the diverse needs of the STIN. Therefore, the SS frameworks for all possible STIN transmission scenarios should be carefully designed, including the uplink and downlink of both TN and NTN. Second, since interference in STIN leads to system performance degradation, mitigating interference becomes the most critical issue. Generally, such mitigation can be implemented by isolating users from the interference source and carefully managing the TN and NTN spectrum [7]. Meanwhile, the TN and NTN spectrum should be allocated appropriately, and the spectrum access scheme should be carefully designed [8]. In addition, since STINs are hierarchical and heterogeneous, the proper selection of performance metrics for system evaluation is also essential to the system, depending on the design objectives.
To address these concerns, we first consider the isolation of STIN users. One of the well-known approaches for this is the utilization of a protection zone for TN, where NTN users within the protection zone are not allowed to access the spectrum shared by TN and NTN. However, the appropriate definition of protection zones still remains a challenging problem. Meanwhile, some studies have been carried out for spectrum allocation and management. Typical work involves spatial SS, dynamic SS, and game-theoretic SS approaches, which are potentially beneficial for STINs. In spatial SS, the objective is to find the geographic areas that are good for NTN access in STINs [9]. In dynamic SS, the objective usually becomes finding the temporal holes in the spectrum resource by treating the NTN as a secondary network [10]. Both spatial SS and dynamic SS can be considered as frequency reuse schemes in STINs. For spectrum allocation and sharing, game models are competitive candidates in finding the optimal scheme by modeling TN and NTN as game players [11]. All of the above topics can be further studied to meet the performance requirements of future STINs.
@@ II-A NTN Downlink and TN Downlink SS
When a satellite transmits signals to an NTN user and a BS transmits signals to a TN user, they can access the shared spectrum, as shown in Fig. 2(1). For example, this framework corresponds to the NTN satellite n256 frequency band and new radio (NR) n65 and n66 frequency bands in the International Telecommunication Union Radiocommunication Sector (ITU-R) and the 3rd Generation Partnership Project (3GPP) [12]. It’s worth noting that SS methods not only apply to the current spectrum regulations but also facilitate future spectrum utilization and standards development. In this case, TN users can be regarded as primary users and NTN users as secondary users. Designing a three-dimensional (3D) protection zone for each TN user is essential. The protection zone can be either spherical or tapered. Satellites outside this protection zone may utilize the shared spectrum while those within continue to use the reserved spectrum. This is because sharing spectrum between TN and NTN will not only enhance the communication performance of TN users through the expanded bandwidth but also result in interference. The protection zone is mainly set to protect the quality of service (QoS) of TN users. Theoretically, NTN users can also be affected by interference leakage from BSs, considering their beam directionality on the antenna surface in the horizontal plane. However, near-vertical transmissions from satellites can hardly be influenced. Moreover, satellites can mitigate mutual interference by constraining spectrum access within a specified distance range.
@@ II-B NTN Downlink and TN Uplink SS
Consider a scenario involving sharing a reverse link spectrum in which an NTN user receives signals from a satellite. As depicted in Fig. 2(2), both the satellite and the NTN user can access the shared spectrum through TN uplink communications. In this framework, satellite spectrum sensing can be performed to detect the strength of TN uplink signals to reduce mutual interference. If the uplink signal exhibits a low level of strength, the NTN downlink transmission can be operated to prevent significant mutual interference. To be more specific, due to the directionality of satellite downlink signals, the corresponding interference to TN uplink transmissions on the near-horizontal plane is relatively small. In addition, owing to the extended propagation distance and limited antenna gain associated with the satellite’s side lobes, TN uplink users located within the reception range of these side lobes exert minimal interference on the satellites’ spectrum sensing capabilities. In particular, spectrum sensing is conducted because satellites and NTN users can neither determine TN users transmitting in the current time slot nor predict aggregated interference.
@@ II-C NTN Uplink and TN Downlink SS
When an NTN user transmits signals to a satellite, they can opportunistically access the shared spectrum with TN downlink communications, as shown in Fig. 2(3). For example, this framework corresponds to the NTN satellite n256 frequency band and NR n2, n25, and n70 frequency bands in ITU-R. This framework is feasible due to a couple of factors. First, the BS is tilted downward for TN downlink communications and employs beamforming techniques to serve TN users. Because of the BS’s antenna angle and beamforming, interference leakage from the BS to the satellite is minimal. Second, during NTN uplink communications, the NTN user and the satellite can collaboratively perform spectrum sensing. This allows them to opportunistically access the spectrum while minimizing interference to and from TN users. It is important to note that, based on spectrum sensing conducted by NTN users, the aggregated interference power generated by NTN users is relatively lower than the desired signal power of the TN downlink communication. This is due to the high transmit power associated with TN BS, which also employs beamforming techniques.
@@ II-D NTN Uplink and TN Uplink SS
In the ITU-R, NTN satellite n256 frequency band and NR n65 frequency band overlap. As shown in Fig. 2(4), allowing NTN uplink communications to share the spectrum with TN uplink communications is facing many problems, which makes it necessary for NTN to use the n256 band in areas where TN does not use the n65 band. First, numerous TN uplink users operate within a satellite’s broad coverage area, and the distances between these users and NTN uplink users can be significant. This results in limited detection capabilities, making it difficult for either type of user to perceive the other effectively. Second, TN users typically have a limited number of antennas, leading to considerable interference leakage from TN users in the vertical direction. This interference can overwhelm the desired signals from NTN uplink communications. Third, NTN users are likely to experience a high rate of missed detections regarding the presence of TN users. Consequently, this inability to accurately sense TN users can lead to significant interference from NTN uplink transmissions to TN communications. As a result, TN users’ communication performance often falls short of expectations. Sharing the spectrum between these two co-directional uplink links is unfavorable from the perspective of interference and sensing capabilities in many situations. However, this spectrum-sharing framework may be viable with appropriate spectrum-aware strategies when the number of TN uplink users is relatively small.
@@ III-A Protection Zone-Based Approach
Method: Protection zones act as an auxiliary approach in SS as they define some geographical area in which users cannot access the spectrum of the primary system. SS techniques enable STIN users to sense and access the spectrum occupied by primary users to improve the spectrum efficiency. However, primary system interference cannot be ignored when users are close to interference sources. As protection zones are usually located around primary users, they can isolate interference for the users inside them, as shown in Fig. 3(a). While the users are outside the protection zones, they are free to access the spectrum.
Applications: With protection zones and knowledge of the user location in STINs, the SS technique enables intelligent monitoring of spatial spectrum holes and shares the spectrum outside the protection zone.
Difficulties and Solutions: Since the NTN is mobile and provides flexible spectrum access for users, finding the boundary for the protection zone becomes a challenging issue. The dynamics of STIN architectures also have significant impacts on the protection zones. Both challenges require statistical modeling and analysis based on user distributions.
@@ III-B Spatial Spectrum Sensing (SSS)-Based Approach
Applications: When SSS is implemented in STINs, NTN-related devices act as secondary users to monitor the busy status of the shared channel. The SSS-based approach suits the NTN downlink and TN uplink SS framework. Acquiring the locations of currently transmitting TN users at NTN satellites is difficult in practical engineering. By selecting an appropriate energy detection threshold, we can manage the interference generated by NTN toward TN.
Difficulties and Solutions: The primary technical challenge of this method is accurately determining the probabilities of false alarms and missed detections. It is crucial to control these probabilities following the Neyman-Pearson criterion. Possible solutions include calculating probability distributions based on the statistical properties of the interference under various events or using machine learning techniques to fit these probabilities.
@@ III-C Joint Spatial Spectrum Sensing (JSSS)-Based Approach
Applications: This approach is particularly suitable for the NTN uplink and TN downlink SS framework. The reason is that the satellite, acting as the NTN receiver, may sometimes experience significant interference from the TN when the number of ground base stations is high in specific areas. Thus, the vertical interference leakage could not be ignored. As a result, the SSS performed by the satellite protects its communication from the interference generated by the TN. At the same time, the SSS done by the NTN user helps manage the interference produced by the NTN towards the TN.
Difficulties and Solutions: Since there are two different types of network nodes in JSSS, collaborating to decide whether SS can be accomplished has become a significant challenge. The energy detection thresholds of the satellite and the NTN user can be jointly optimized under the Neyman-Pearson criterion.
@@ III-D Dynamic Spectrum Access-Based Approach
Applications: Since NTNs have high mobility, STIN users can adopt DSA strategies to access the NTN spectrum by integrating artificial intelligence (AI) tools. Similarly, NTNs can sense the spectrum licensed to TN users and adopt DSA to provide services for NTN users. Hence, the DSA protocol is essential for the DSA-based approach in SS and has been designed to guarantee the performance of STINs in existing studies.
Difficulties and Solutions: Since the NTN nodes need to sense and detect spectrum holes at various geolocations over time to provide wireless access, machine learning techniques incorporating dedicated geometric models are necessary for DSA in STINs. Meanwhile, since NTN nodes are independent, distributed NTN sensing schemes should be designed as well.
@@ III-E Game Theory-Based Approach
Method: Game theory is a powerful theoretical approach that models the interacting users as rational players aiming to maximize their own benefits by selling and buying spectrum resources, as shown in Fig. 3(e). In STINs, users tend to make their decisions and compete with each other to access the shared spectrum, and ultimately decide the optimal SS pattern by obtaining the game Nash equilibrium.
Applications: In STINs, the spectrum allocation and access scheme of the TN/NTN nodes and the users can be described by adopting precise game models, and the optimal spectrum utilization pattern can be obtained. This is particularly suitable for NTN downlink scenarios, as potential users may access the shared spectrum when deciding their access mode.
Difficulties and Solutions: Since STINs are heterogeneous with multiple hierarchies, the internal coordination of various users and network devices in SS often exhibits considerably high complexity. In this sense, the selection of appropriate game theory approach is the key issue in modeling the sophisticated interaction behaviors of STINs.

### v3

@@ Abstract
With the construction of low-earth orbit (LEO) satellite constellations, ubiquitous connectivity has been achieved. Terrestrial networks (TNs), such as cellular networks, are mainly deployed in specific urban areas and use licensed spectrum. However, in remote areas where terrestrial infrastructure is sparse, licensed spectrum bands are often underutilized. To accommodate the increasing communication needs, non-terrestrial networks (NTNs) can opportunistically access this idle spectrum to improve spectrum efficiency via spectrum sharing (SS). Therefore, bringing NTNs to a shared spectrum with TNs can improve network capacity under reasonable interference management. In satellite-terrestrial integrated networks (STINs), the comprehensive coverage of a satellite and the unbalanced communication resources of STINs make it challenging to manage mutual interference between NTN and TN effectively. This article presents the fundamentals and prospects of SS in STINs by introducing four SS frameworks, their potential application scenarios, and technical challenges. Furthermore, advanced SS approaches related to interference management in STINs and performance metrics of SS in STINs are introduced. Moreover, a preliminary performance evaluation showcases the potential for sharing the spectrum between NTN and TN. Finally, future research opportunities for SS in STINs are discussed.
@@ I Introduction
Due to the growing need for broader and more reliable wireless network coverage and data transmission, non-terrestrial networks (NTNs) have emerged as a promising solution. In urban areas with dense users, the bandwidth shortage is a primary challenge, while in remote areas, the main problem is limited coverage. However, in such remote regions, a large portion of licensed terrestrial spectrum remains underutilized due to the sparse deployment of base stations [1]. By allowing NTNs to opportunistically access these idle spectrum bands, spectrum sharing (SS) in STINs can significantly improve overall spectrum utilization without increasing allocation demands. Spectrum integration in satellite-terrestrial integrated networks (STINs) is a necessary step to improve the utilization efficiency of scarce spectrum resources. However, the coexistence of NTNs and TNs can be challenging, particularly in terms of spectrum utilization and interference management. When NTN nodes, such as low-earth orbit (LEO) satellites, are incorporated into existing TNs via specific frequency bands, interference patterns can become more complex due to the different spatial propagation characteristics of signals. For example, the use of the Ka band or the Ku band for satellites may cause severe co-channel interference in existing communication systems. Moreover, within a fixed spectrum allocation, it is possible that while certain portions of the spectrum in NTNs are fully occupied, other segments within TNs may remain unutilized. The imbalance will lead to under-utilization of available spectrum resources and a significant hindrance to substantial improvements in system performance.
@@
In recent years, the rapid growth of NTNs has paved the way toward realizing STINs. STINs, a promising candidate for future sixth-generation (6G) wireless communications, incorporate LEO nodes into existing TN to take advantage of NTN and meet diverse communication requirements [2]. By using STINs, the coverage, flexibility, and spectrum efficiency of the wireless network can be significantly improved, as terrestrial users can access various TN and NTN nodes from vast geographical areas. Inevitably, the licensed spectrum resource will become scarcer in STIN as more NTN nodes are introduced. This encourages us to manage the licensed spectrum when TN and NTN co-exist in the STIN [3]. Meanwhile, the licensed spectrum to the static TN nodes cannot always be utilized efficiently, as spatial spectrum holes cannot be eliminated within the geographic area of interest. Hence, mobile NTN users can admit access to users within the spectrum holes and promote spectrum efficiency [4]. In existing work, one of the promising technologies for TN and NTN coexistence is SS, which allows the NTN to share the spectrum pre-allocated to the TN. SS allows the NTN to work in the same frequency band when the TN users are active, as long as its interference to the TN can be tolerated [4, 5]. In this way, the STIN can provide high-quality and seamless wireless services at a slight cost of performance loss under the practical restrictions of limited spectrum resources.
@@
With the distinct advantages mentioned above, STINs can outperform conventional TNs in various practical fields. As the key technique in STINs, SS becomes the main bottleneck in network performance promotion and faces several challenges [6]. First, a universal SS architecture cannot always meet the diverse needs of the STIN. Therefore, the SS frameworks for all possible STIN transmission scenarios should be carefully designed, including the uplink and downlink of both TN and NTN. Second, since interference in STIN leads to system performance degradation, mitigating interference becomes the most critical issue. Generally, such mitigation can be implemented by isolating users from the interference source and carefully managing the TN and NTN spectrum [7]. Meanwhile, the TN and NTN spectrum should be allocated appropriately, and the spectrum access scheme should be carefully designed [8]. In addition, since STINs are hierarchical and heterogeneous, the proper selection of performance metrics for system evaluation is also crucial, depending on the design objectives.
Apart from the above concerns, one factor that has a significant impact on SS is the framework in which the TN and NTN nodes operate, whether in uplink or downlink mode. In various frameworks, the mutual interaction between TN and NTN networks, as well as their transmission characteristics, may vary significantly, which are primary factors in designing proper SS strategies. Based on the SS frameworks, several state-of-the-art SS techniques are considered to address these concerns. For instance, a protection zone can be established to prevent NTN users within that zone from accessing the spectrum shared by TN and NTN. However, the defining protection zones remain a significant challenge. In addition, spectrum sensing is another crucial area for SS and is beneficial for STINs. Representative techniques, such as spatial spectrum sensing (SSS) [9] and dynamic spectrum sensing [10], enable us to detect and reuse spatial and temporal spectrum holes within the STIN spectrum, respectively. In terms of spectrum allocation, game-theoretic models are competitive candidates, as they can represent TN and NTN as competing players [11]. All of these topics warrant further study to address the performance requirements of future STINs.
@@ II-A NTN Downlink and TN Downlink SS
When a satellite transmits signals to an NTN user and a BS transmits signals to a TN user, they can access the shared spectrum, as shown in Fig. 2(1). For example, this framework corresponds to the NTN satellite n256 frequency band and new radio (NR) n65 and n66 frequency bands in the International Telecommunication Union Radiocommunication Sector (ITU-R) and the 3rd Generation Partnership Project (3GPP) [12]. It’s worth noting that SS methods not only apply to the current spectrum regulations but also facilitate future spectrum utilization and standards development. In this case, TN users can be regarded as primary users and NTN users as secondary users. Designing a three-dimensional (3D) protection zone for each TN user is essential. The protection zone can be either spherical or tapered. Satellites outside this protection zone may utilize the shared spectrum while those within continue to use the reserved spectrum. This is because sharing spectrum between TN and NTN will not only enhance the communication performance of TN users through the expanded bandwidth but also result in interference. The protection zone is primarily established to maintain the quality of service (QoS) for TN users. Theoretically, NTN users can also be affected by interference leakage from BSs, considering their beam directionality on the antenna surface in the horizontal plane. However, near-vertical transmissions from satellites can hardly be influenced. Moreover, satellites can mitigate mutual interference by constraining spectrum access within a specified distance range.
@@ II-B NTN Downlink and TN Uplink SS
Consider a scenario involving the sharing of a reverse link spectrum, in which an NTN user receives signals from a satellite. As depicted in Fig. 2(2), both the satellite and the NTN user can access the shared spectrum through TN uplink communications. In this framework, satellite spectrum sensing can be performed to detect the strength of TN uplink signals, thereby reducing mutual interference. If the uplink signal exhibits a low level of strength, the NTN downlink transmission can be operated to prevent significant mutual interference. To be more specific, due to the directionality of satellite downlink signals, the corresponding interference to TN uplink transmissions on the near-horizontal plane is relatively small. In addition, owing to the extended propagation distance and limited antenna gain associated with the satellite’s side lobes, TN uplink users located within the reception range of these side lobes exert minimal interference on the satellite’s spectrum sensing capabilities. In particular, spectrum sensing is conducted because satellites and NTN users cannot determine TN users transmitting in the current time slot or predict aggregated interference.
@@ II-C NTN Uplink and TN Downlink SS
When an NTN user transmits signals to a satellite, they can opportunistically access the shared spectrum with TN downlink communications, as shown in Fig. 2(3). For example, this framework corresponds to the NTN satellite n256 frequency band and NR n2, n25, and n70 frequency bands in ITU-R. This framework is feasible due to several factors. First, the BS is tilted downward for TN downlink communications and employs beamforming techniques to serve TN users. Due to the BS’s antenna angle and beamforming, interference leakage from the BS to the satellite is minimized. Second, during NTN uplink communications, the NTN user and the satellite can collaboratively perform spectrum sensing. This allows them to opportunistically access the spectrum while minimizing interference to and from TN users. It is important to note that, based on spectrum sensing conducted by NTN users, the aggregated interference power generated by NTN users is relatively lower than the desired signal power of the TN downlink communication. This is due to the high transmit power associated with TN BS, which also employs beamforming techniques.
@@ II-D NTN Uplink and TN Uplink SS
In the ITU-R, the NTN n256 frequency band and the NR n65 frequency band overlap. As shown in Fig. 2(4), allowing NTN uplink communications to share the spectrum with TN uplink communications poses several problems, necessitating that NTN use the n256 band in areas where TN does not use the n65 band. First, numerous TN uplink users operate within a satellite’s broad coverage area, and the distances between these users and NTN uplink users can be significant. This results in limited detection capabilities, making it difficult for either type of user to perceive the other effectively. Second, TN users typically have a limited number of antennas, leading to considerable interference leakage from TN users in the vertical direction. This interference can overwhelm the desired signals from the NTN uplink communications. Third, NTN users are likely to experience a high rate of missed detections regarding the presence of TN users. Consequently, this inability to accurately sense TN users can lead to significant interference from NTN uplink transmissions to TN communications. As a result, the communication performance of TN users often falls short of expectations. Sharing the spectrum between these two co-directional uplink links is unfavorable from the perspective of interference and sensing capabilities in many situations. However, this spectrum-sharing framework may be viable with appropriate spectrum-aware strategies when the number of TN uplink users is relatively small.
@@ III-A Protection Zone-Based Approach
Method: Protection zones serve as an auxiliary approach in SS, as they define a geographical area in which users cannot access the spectrum of the primary system. SS techniques enable STIN users to sense and access the spectrum occupied by primary users to improve the spectrum efficiency. However, primary system interference cannot be ignored when users are close to interference sources. As protection zones are usually located around primary users, they can isolate interference for the users inside them, as shown in Fig. 3(a). While the users are outside the protection zones, they are free to access the spectrum.
Applications: With protection zones and user locations, the SS technique enables us to guarantee the key TN device performance and protect them from outage and malicious attacks. SS also enables intelligent monitoring of the spatial spectrum and spectrum sharing outside the protection zone.
Difficulties and Solutions: Since NTN is mobile and provides flexible spectrum access, finding the protection zone boundary becomes the major challenge. The dynamics of STIN architectures also impose another challenge on the protection zone. Both challenges require signal statistical modeling and analysis based on user distributions.
@@ III-B Spatial Spectrum Sensing (SSS)-Based Approach
Applications: When SSS is implemented in STINs, NTN-related devices act as secondary users to monitor the busy status of the shared channel. By selecting an appropriate energy detection threshold, we can achieve efficient interference management from NTN toward TN. Meanwhile, we can provide spectrum information for terrestrial devices and thus support better service coverage.
Difficulties and Solutions: The major challenge of this method is precisely modeling signals and interference, which determines the false alarm and missed detection probabilities. It is crucial to control these probabilities following the Neyman-Pearson criterion. Possible solutions include calculating probability distributions based on interference statistical properties under various events or using machine learning techniques to fit these probabilities.
@@ III-C Joint Spatial Spectrum Sensing (JSSS)-Based Approach
Applications: This approach is suitable for NTN uplink and TN downlink SS framework, as the satellite may experience significant interference from the TN when the number of ground base stations is high in specific areas. As a result, the SSS protects NTN communication from TN interference. Meanwhile, the SSS done by the NTN user helps manage the interference produced by the NTN towards the TN. In addition, SSS data analysis can help us find out abnormal historical data and better utilize the system spectrum resource.
Difficulties and Solutions: Since there are two different types of network nodes in JSSS, the joint decision whether SS can be accomplished has become the major challenge. To address this issue, the energy detection thresholds of the satellite and the NTN user can be jointly optimized under the Neyman-Pearson criterion.
@@ III-D Dynamic Spectrum Access-Based Approach
Applications: Since NTNs have high mobility, STIN users can adopt DSA strategies to access the NTN spectrum. Similarly, NTNs can sense the spectrum licensed to TN users and adopt DSA to provide services for NTN users. Hence, flexible and robust networking in STINs can be achieved by adopting DSA. Meanwhile, DSA also enables flexible network access in remote areas where TN is no longer available.
Difficulties and Solutions: The sensing and decision mechanisms in DSA pose major challenges. To support these functionalities, the DSA protocol is also essential. Since NTN nodes need to detect spectrum holes at various geolocations over time, machine learning techniques incorporating dedicated geometric models are necessary for DSA in STINs. Meanwhile, since the NTN nodes are independent, the corresponding sensing schemes and protocols should also be designed.
@@ III-E Game Theory-Based Approach
Method: Game theory is a powerful theoretical approach that models the interacting users as rational players aiming to maximize their benefits by selling and buying spectrum resources, as shown in Fig. 3(e). In STINs, users tend to make their decisions and compete with each other to access the shared spectrum, ultimately deciding on the optimal SS pattern by reaching the game’s Nash equilibrium.
Applications: In STINs, the adoption of game theory models allows us to coordinate the interactions among TN and NTN nodes under potential spectrum conflict, thus maintaining TN and NTN performance as well as maximizing spectrum utilization. Game models can also be merged with DSA to decide the spectrum access patterns. This is suitable for NTN downlink and low-altitude communications, as the users can access the shared spectrum when deciding their access mode.
Difficulties and Solutions: Since STINs are hierarchical and heterogeneous, the internal coordination of various devices in SS often exhibits considerable complexity. To address the problem, selecting the appropriate network model and game model is a key issue in spectrum sensing, access, and management of STINs.